# Robust Online Conformal Prediction under Uniform Label Noise

### v2

@@ Abstract
Conformal prediction is an emerging technique for uncertainty quantification that constructs prediction sets guaranteed to contain the true label with a predefined probability. Recent work develops online conformal prediction methods that adaptively construct prediction sets to accommodate distribution shifts. However, existing algorithms typically assume perfect label accuracy which rarely holds in practice. In this work, we investigate the robustness of online conformal prediction under uniform label noise with a known noise rate, in both constant and dynamic learning rate schedules. We show that label noise causes a persistent gap between the actual mis-coverage rate and the desired rate $\alpha$, leading to either overestimated or underestimated coverage guarantees. To address this issue, we propose Noise Robust Online Conformal Prediction (dubbed NR-OCP) by updating the threshold with a novel robust pinball loss, which provides an unbiased estimate of clean pinball loss without requiring ground-truth labels. Our theoretical analysis shows that NR-OCP eliminates the coverage gap in both constant and dynamic learning rate schedules, achieving a convergence rate of $\mathcal{O}(T^{-1/2})$ for both empirical and expected coverage errors under uniform label noise. Extensive experiments demonstrate the effectiveness of our method by achieving both precise coverage and improved efficiency.
Machine Learning, ICML
@@ 1 Introduction
Machine learning techniques are revolutionizing decision-making in high-stakes domains, such as autonomous driving (Bojarski et al., 2016) and medical diagnostics (Caruana et al., 2015). It is crucial to ensure the reliability of model predictions in these contexts, as wrong predictions can result in serious consequences. While various techniques have been developed for uncertainty estimation, including confidence calibration (Guo et al., 2017) and Bayesian neural networks (Smith, 2013), they typically lack rigorous theoretical guarantees. Conformal prediction addresses this limitation by establishing a systematic framework to construct prediction sets with provable coverage guarantee (Vovk et al., 2005; Shafer & Vovk, 2008; Balasubramanian et al., 2014; Angelopoulos et al., 2023). Notably, this framework requires no parametric assumptions about the data distribution and can be applied to any black-box predictor, which makes it a powerful technique for uncertainty quantification.
@@
Recent research extends conformal prediction to online setting where the data arrives in a sequential order (Gibbs & Candes, 2021; Feldman et al., 2023; Bhatnagar et al., 2023; Angelopoulos et al., 2024; Gibbs & Candès, 2024; Hajihashemi & Shen, 2024). These methods provably achieve the desired coverage property under arbitrary distributional changes. However, previous studies typically assume perfect label accuracy, an assumption that seldom holds true in practice due to the common occurrence of noisy labels in online learning (Ben-David et al., 2009; Natarajan et al., 2013; Wu et al., 2024). Recent work (Einbinder et al., 2024) proves that online conformal prediction can achieve a conservative coverage guarantee under uniform label noise, leading to unnecessarily large prediction sets. However, their analysis relies on a strong distributional assumption of non-conformity score and cannot quantify the specific deviation of coverage guarantees. These limitations motivate us to establish a general theoretical framework for this problem and develop a noise-robust algorithm that maintains precise coverage guarantees while producing small prediction sets.
@@
In this work, we analyze online conformal prediction under uniform label noise with a known noise rate, in both constant (Gibbs & Candes, 2021) and dynamic (Angelopoulos et al., 2024) learning rate schedule. Notably, our theoretical results are independent of the distributional assumption made in the previous work (Einbinder et al., 2024). In particular, we demonstrate that label noise causes a persistent gap between the actual mis-coverage rate and the desired rate $\alpha$, with higher noise rates resulting in larger gaps. This gap can lead to either overestimated or underestimated coverage guarantees, which depend on the size of the prediction sets.
@@
To address this challenge, we propose Noise-Robust Online Conformal Prediction (dubbed NR-OCP) by updating the threshold with a robust pinball loss, which provides an unbiased estimate of clean pinball loss value without requiring ground-truth labels. Specifically, we construct the robust pinball loss as a weighted combination of the pinball loss with noisy scores and the pinball loss with scores of all classes. We prove that this loss is equivalent to the pinball loss under clean labels in expectation. Our theoretical analysis shows that NR-OCP eliminates the coverage gap, caused by the label noise, in both constant and dynamic learning rate schedules, achieving a convergence rate of $\mathcal{O}(T^{-1/2})$ for both empirical and expected coverage errors (i.e., absolute deviation of the empirical and expected mis-coverage rate from the target level $\alpha$) under the uniform label noise.
@@
To verify the effectiveness of our method, we conduct extensive experiments on CIFAR-100 (Krizhevsky, 2009) and ImageNet (Deng et al., 2009) with synthetic uniform label noise. Empirical results show that NR-OCP consistently achieves the desired $1-\alpha$ long-run coverage rate, while the standard online conformal prediction updated with noisy labels (baseline) shows a significant coverage gap. For example, using DenseNet121 on CIFAR-100 dataset with noise rate $\epsilon=0.05$, error rate $\alpha=0.1$, and constant learning rate, the baseline method (Gibbs & Candes, 2021) significantly deviates from the target coverage level of $0.9$ using the LAC score, exhibiting a coverage gap of 3.178% and an average prediction set size of 25.88. In contrast, NR-OCP achieves a negligible coverage gap of 0.156% and a smaller average prediction set size of 15.20. Notably, our method demonstrates superior performance over the baseline across different settings, including various model architectures, error rates, noise rates, and non-conformity scores.
@@
  • We present a general theoretical framework for analyzing the impact of uniform label noise on online conformal prediction under constant (Proposition 3.1) and dynamic (Proposition 4.1) learning rates, respectively. Our theoretical results are free from the distributional assumption in prior work (Einbinder et al., 2024).
  • We introduce NR-OCP, a novel method of online conformal prediction that is robust to uniform label noise. Our method includes a robust pinball loss that provides an unbiased estimate of clean pinball loss without requiring access to clean labels.
  • We prove that NR-OCP eliminates the coverage gap in online conformal prediction with both constant (Propositions 3.3 and 3.4) and dynamic learning rate (Propositions 4.3 and 4.4), achieving convergence rates of $\mathcal{O}(T^{-1/2})$ for both empirical and expected coverage errors under uniform label noise.
@@ 2 Preliminary
We study the problem of generating prediction sets in online classification where the data arrives in a sequential order (Gibbs & Candes, 2021; Angelopoulos et al., 2024). Formally, we consider a sequence of data points $(X_{t},Y_{t})$, $t\in\mathbb{N}^{+}$, which are sampled from a joint distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ over the input space $\mathcal{X}\subset\mathbb{R}^{d}$, and the label space $\mathcal{Y}=\{1,\ldots,K\}$. In online conformal prediction, the goal is to construct prediction sets $\mathcal{C}_{t}(X_{t})$, $t\in\mathbb{N}^{+}$, that provides precise coverage guarantee: $\lim_{T\to+\infty}\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}=\alpha$, where $\alpha\in(0,1)$ denotes a user-specified error rate.
@@
$\mathcal{C}_{t}(X_{t})=\left\{y\in\mathcal{Y}\,:\,\mathcal{S}(X_{t},y)\leq\hat{\tau}_{t}\right\}.$
where $\hat{\tau}_{t}$ is a data-driven threshold, and $\mathcal{S}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ denotes a non-conformity score function that measures the deviation between a data sample and the training data. For example, given a pre-trained classifier $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$, the LAC score (Sadinle et al., 2019) is defined as $\mathcal{S}(X,Y)=1-\hat{\pi}_{Y}(X)$, where $\hat{\pi}_{Y}(X)=\sigma_{Y}(f(X))$ denotes the softmax probability of instance $X$ for class $Y$, and $\sigma$ is the softmax function. For notation shorthand, we use $S_{t}$ to denote the random variable $\mathcal{S}(X_{t},Y_{t})$ and use $S_{t,y}$ to denote $\mathcal{S}(X_{t},y)$ for a given class $y\in\mathcal{Y}$. Following previous work (Kiyani et al., 2024; Angelopoulos et al., 2024), we will assume that the non-conformity score function is bounded, and the threshold is specifically initialized:
@@
After the prediction, we can observe the corresponding label $Y_{t}$ for $X_{t}$. Given a desired mis-coverage rate $\alpha$, online conformal prediction (Gibbs & Candes, 2021; Angelopoulos et al., 2024) updates the threshold $\hat{\tau}_{t}$ with pinball loss:
@@
where $\tau$ denotes a threshold and $s$ is a non-conformity score. Then, the threshold is updated via online gradient descent:
$\hat{\tau}_{t+1}$ $=\hat{\tau}_{t}-\eta_{t}\cdot\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S_{t})$ (1)
$=\hat{\tau}_{t}+\eta_{t}\cdot(\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}-\alpha)$
where $\nabla_{\tau}l_{1-\alpha}(\tau,s)$ denotes the gradient of pinball loss w.r.t the threshold $\tau$, and $\eta_{t}>0$ is the learning rate. The optimization will increase the threshold if the prediction set $\mathcal{C}_{t}(X_{t})$ fails to encompass the label $Y_{t}$, resulting in more conservative predictions in future instances (and vice versa).
We use the empirical coverage error and the expected coverage error to evaluate the coverage performance. The empirical coverage error measures the absolute deviation of the mis-coverage rate from the target level $\alpha$:
$\mathrm{EmErr(T)}=\left|\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}-\alpha\right|,$
while the expected coverage error quantifies the absolute deviation in expectation:
$\mathrm{ExErr(T)}=\left|\frac{1}{T}\sum_{t=1}^{T}\mathbb{P}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}-\alpha\right|.$
In addition, online learning theory (see Theorem 2.13 of Orabona (2019)) established that standard online conformal prediction achieves a regret bound:
$\mathrm{Reg}(T)$ $=\sum_{t=1}^{T}l_{1-\alpha}(\hat{\tau}_{t},S_{t})-\min_{\tau}\sum_{t=1}^{T}l_{1-\alpha}(\tau,S_{t})$
$=\mathcal{O}(T^{-\frac{1}{2}}),$
with optimally chosen $\eta_{t}$, which serves as a helpful measure alongside coverage (Bhatnagar et al., 2023). We provide a regret analysis for our method in Appendix B.1.
Label noise.
In this paper, we focus on the issue of noisy labels in online learning, a common occurrence in the real world. This is primarily due to the dynamic nature of real-time data streams and the potential for human error or sensor malfunctions during label collection. Let $(X_{t},\tilde{Y}_{t})$ be the data sequence with label noise, and $\tilde{S}_{t}=\mathcal{S}(X_{t},\tilde{Y}_{t})$ be the noisy non-conformity score. In this work, we focus on the setting of uniform label noise  (Einbinder et al., 2024; Penso & Goldberger, 2024), i.e., the correct label is replaced by a label that is randomly sampled from the $K$ classes with a fixed probability $\epsilon\in(0,1)$:
@@
where $U$ is uniformly distributed over $[0,1]$, and $\bar{Y}$ is uniformly sampled from the set of classes $\mathcal{Y}$. We assume the probability $\epsilon$ (i.e., the noise rate) is known, in alignment with prior works (Einbinder et al., 2024; Penso & Goldberger, 2024; Sesia et al., 2024). This assumption is practical as the noise rate can be estimated from historical data (Liu & Tao, 2015; Yu et al., 2018; Wei et al., 2020).
Recent work (Einbinder et al., 2024) investigates the noise robustness of online conformal prediction under uniform label noise, with a strong distributional assumption. Their analysis demonstrates that noisy labels will lead to a conservative long-run coverage guarantee, with the assumption that noisy score distribution stochastically dominates the clean score distribution, i.e., $\mathbb{P}\left\{\tilde{S}\leq s\right\}\leq\mathbb{P}\left\{S\leq s\right\},\forall s\in\mathbb{R}$. The distributional assumption is strong so that it only ensures valid coverage under limited cases. Moreover, their analysis fails to quantify the specific deviation of coverage guarantees. These limitations motivate us to establish a general theoretical framework for this problem and develop a noise-robust algorithm for online conformal prediction. We proceed by presenting the theoretical results on both constant (Gibbs & Candes, 2021) and dynamic (Angelopoulos et al., 2024) learning rate schedules, respectively.
3 Results for constant learning rate
3.1 The impact of label noise on prediction sets
In this section, we theoretically analyze the impacts of uniform label noise on online conformal prediction with a constant learning rate, i.e., $\eta_{t}\equiv\eta$, for all $t\in\mathbb{N}^{+}$. In this case, the threshold $\hat{\tau}_{t}$ is updated by

### v3

# Exploring the Noise Robustness of Online Conformal Prediction

@@ Abstract
Conformal prediction is an emerging technique for uncertainty quantification that constructs prediction sets guaranteed to contain the true label with a predefined probability. Recent work develops online conformal prediction methods that adaptively construct prediction sets to accommodate distribution shifts. However, existing algorithms typically assume perfect label accuracy which rarely holds in practice. In this work, we investigate the robustness of online conformal prediction under uniform label noise with a known noise rate. We show that label noise causes a persistent gap between the actual mis-coverage rate and the desired rate $\alpha$, leading to either overestimated or underestimated coverage guarantees. To address this issue, we propose a novel loss function robust pinball loss, which provides an unbiased estimate of clean pinball loss without requiring ground-truth labels. Theoretically, we demonstrate that robust pinball loss enables online conformal prediction to eliminate the coverage gap under uniform label noise, achieving a convergence rate of $\mathcal{O}(T^{-1/2})$ for both empirical and expected coverage errors (i.e., absolute deviation of the empirical and expected mis-coverage rate from the target level $\alpha$). This loss offers a general solution to the uniform label noise, and is complementary to existing online conformal prediction methods. Extensive experiments demonstrate that robust pinball loss enhances the noise robustness of various online conformal prediction methods by achieving a precise coverage guarantee and improved efficiency.
@@ 1 Introduction
Machine learning techniques are revolutionizing decision-making in high-stakes domains, such as autonomous driving (bojarski2016end, ) and medical diagnostics (caruana2015intelligible, ). It is crucial to ensure the reliability of model predictions in these contexts, as wrong predictions can result in serious consequences. While various techniques have been developed for uncertainty estimation, including confidence calibration (guo2017calibration, ) and Bayesian neural networks (smith2013uncertainty, ), they typically lack rigorous theoretical guarantees. Conformal prediction addresses this limitation by establishing a systematic framework to construct prediction sets with provable coverage guarantee (vovk2005algorithmic, ; shafer2008tutorial, ; balasubramanian2014conformal, ; angelopoulos2023conformal, ). Notably, this framework requires no parametric assumptions about the data distribution and can be applied to any black-box predictor, which makes it a powerful technique for uncertainty quantification.
@@
Recent research extends conformal prediction to online setting where the data arrives in a sequential order (gibbs2021adaptive, ; feldman2022achieving, ; bhatnagar2023improved, ; DBLP:conf/icml/AngelopoulosBB24, ; gibbs2024conformal, ; hajihashemimulti, ). These methods provably achieve the desired coverage property under arbitrary distributional changes. However, previous studies typically assume perfect label accuracy, an assumption that seldom holds true in practice due to the common occurrence of noisy labels in online learning (ben2009agnostic, ; natarajan2013learning, ; wuinformation, ). Recent work (einbinder2024label, ) proves that online conformal prediction can achieve a conservative coverage guarantee under uniform label noise, leading to unnecessarily large prediction sets. However, their analysis relies on a strong distributional assumption of non-conformity score and cannot quantify the specific deviation of coverage guarantees. These limitations motivate us to establish a general theoretical framework for this problem and develop a noise-robust algorithm that maintains precise coverage guarantees while producing small prediction sets.
@@
In this work, we present a general theoretical framework for analyzing how uniform label noise affects the performance of standard online conformal prediction, i.e., adaptive conformal inference (dubbed ACI) (gibbs2021adaptive, ). Notably, our theoretical results are independent of the distributional assumptions on the non-conformity scores made in the previous work (einbinder2024label, ). In particular, we demonstrate that label noise causes a persistent gap between the actual mis-coverage rate and the desired rate $\alpha$, with higher noise rates resulting in larger gaps. This gap can lead to either overestimated or underestimated coverage guarantees, which depend on the size of the prediction sets (see Proposition 3.1).
@@
To address this challenge, we propose a novel loss function robust pinball loss, which provides an unbiased estimate of clean pinball loss value without requiring access to ground-truth labels. Specifically, we construct the robust pinball loss as a weighted combination of the pinball loss computed with respect to noisy scores and the pinball loss with scores of all classes. We prove that this loss is equivalent to the pinball loss under clean labels in expectation. Theoretically, we demonstrate that robust pinball loss enables ACI to eliminate the coverage gap under uniform label noise. It achieves a convergence rate of $\mathcal{O}(T^{-1/2})$ for both empirical and expected coverage errors (i.e., absolute deviation of the empirical and expected mis-coverage rate from the target level $\alpha$). Notably, our robust pinball loss offers a general solution to the uniform label noise, and is complementary to existing online conformal prediction methods.
@@
To verify the effectiveness of the robust pinball loss, we conduct extensive experiments on CIFAR-100 (krizhevsky2009learning, ) and ImageNet deng2009imagenet with synthetic uniform label noise. In particular, we integrate the proposed loss into ACI with constant (gibbs2021adaptive, ) and dynamic learning rates (DBLP:conf/icml/AngelopoulosBB24, ), and strongly adaptive online conformal prediction (bhatnagar2023improved, ). Empirical results show that the robust pinball loss enhances the noise robustness of online conformal prediction by eliminating the coverage gap caused by the label noise. Thus, these methods achieve both long-run coverage and local coverage rate that are close to the target $1-\alpha$, and improved prediction set efficiency. For example, on ImageNet with error rate $\alpha=0.1$, noise rate $\epsilon=0.15$ and dynamic learning rate, ACI with the standard pinball loss deviates from the target coverage level of $0.9$, exhibiting a coverage gap of 8.372% and an average set size of 171.2. In contrast, ACI equipped with the robust pinball loss achieves a negligible coverage gap of 0.183% and a prediction set size of 13.10. In summary, our method consistently enhances the noise robustness of various online conformal prediction methods by achieving a precise coverage guarantee and improved efficiency.
@@
  • We present a general theoretical framework for analyzing the effect of uniform label noise on the coverage of online conformal prediction. Our theoretical results are independent of the distributional assumptions made in the previous work (einbinder2024label, ).
  • To address the issue of label noise, we propose a novel loss function – robust pinball loss that enhances the noise robustness of online conformal prediction. This loss is complementary to online conformal prediction algorithms and can be seamlessly integrated with these methods.
  • We empirically validate that our method can be applied to various online conformal prediction methods and non-conformity score functions. It is straightforward to implement and does not require sophisticated changes to the framework of online conformal prediction.
@@ 2 Preliminary
We study the problem of generating prediction sets in online classification where the data arrives in a sequential order (gibbs2021adaptive, ; DBLP:conf/icml/AngelopoulosBB24, ). Formally, we consider a sequence of data points $(X_{t},Y_{t})$, $t\in\mathbb{N}^{+}$, which are sampled from a joint distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ over the input space $\mathcal{X}\subset\mathbb{R}^{d}$, and the label space $\mathcal{Y}=\{1,\ldots,K\}$. In online conformal prediction, the goal is to construct prediction sets $\mathcal{C}_{t}(X_{t})$, $t\in\mathbb{N}^{+}$, that provides precise coverage guarantee: $\lim_{T\to+\infty}\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}=\alpha$, where $\alpha\in(0,1)$ denotes a user-specified error rate.
@@
$\mathcal{C}_{t}(X_{t})=\left\{y\in\mathcal{Y}\,:\,\mathcal{S}(X_{t},y)\leq\hat{\tau}_{t}\right\}.$ (1)
where $\hat{\tau}_{t}$ is a data-driven threshold, and $\mathcal{S}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ denotes a non-conformity score function that measures the deviation between a data sample and the training data. For example, given a pre-trained classifier $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$, the LAC score (sadinle2019least, ) is defined as $\mathcal{S}(X,Y)=1-\hat{\pi}_{Y}(X)$, where $\hat{\pi}_{Y}(X)=\sigma_{Y}(f(X))$ denotes the softmax probability of instance $X$ for class $Y$, and $\sigma$ is the softmax function. For notation shorthand, we use $S_{t}$ to denote the random variable $\mathcal{S}(X_{t},Y_{t})$ and use $S_{t,y}$ to denote $\mathcal{S}(X_{t},y)$ for a given class $y\in\mathcal{Y}$. Following previous work (DBLP:conf/icml/AngelopoulosBB24, ; DBLP:conf/icml/KiyaniPH24, ), we will assume that the non-conformity score function is bounded, and the threshold is specifically initialized:
@@
In online conformal prediction, a representative method is adaptive conformal inference (ACI) (gibbs2021adaptive, ), which updates the threshold $\hat{\tau}_{t}$ with pinball loss:
@@
where $\tau$ denotes a threshold and $s$ is a non-conformity score. As the label $Y_{t}$ of $X_{t}$ is observed after model prediction, the threshold is then updated via online gradient descent:
$\hat{\tau}_{t+1}=\hat{\tau}_{t}-\eta\cdot\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S_{t})=\hat{\tau}_{t}+\eta\cdot(\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}-\alpha)$ (2)
where $\nabla_{\tau}l_{1-\alpha}(\tau,s)$ denotes the gradient of pinball loss w.r.t the threshold $\tau$, and $\eta>0$ is the learning rate. The optimization will increase the threshold if the prediction set $\mathcal{C}_{t}(X_{t})$ fails to encompass the label $Y_{t}$, resulting in more conservative predictions in future instances (and vice versa).
We use the empirical coverage error and the expected coverage error to evaluate the coverage performance. The empirical coverage error measures the absolute deviation of the mis-coverage rate from the target level $\alpha$, while the expected coverage error quantifies the absolute deviation in expectation. In particular, for any $T\in\mathbb{N}^{*}$, we define
$\mathrm{EmErr}(T)=\left|\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}-\alpha\right|,\quad\mathrm{ExErr}(T)=\left|\frac{1}{T}\sum_{t=1}^{T}\mathbb{P}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}-\alpha\right|.$
Uniform Label noise.
In this paper, we focus on the issue of noisy labels in online learning, a common occurrence in the real world. This is primarily due to the dynamic nature of real-time data streams and the potential for human error or sensor malfunctions during label collection. Let $(X_{t},\tilde{Y}_{t})$ be the data sequence with label noise, and $\tilde{S}_{t}=\mathcal{S}(X_{t},\tilde{Y}_{t})$ be the noisy non-conformity score. In this work, we focus on the setting of uniform label noise  (einbinder2024label, ; penso2024conformal, ), i.e., the correct label is replaced by a label that is randomly sampled from the $K$ classes with a fixed probability $\epsilon\in(0,1)$:
@@
where $U$ is uniformly distributed over $[0,1]$, and $\bar{Y}$ is uniformly sampled from the set of classes $\mathcal{Y}$. We assume the probability $\epsilon$ (i.e., the noise rate) is known, in alignment with prior works (einbinder2024label, ; penso2024conformal, ; sesia2024adaptive, ). This assumption is practical as the noise rate can be estimated from historical data liu2015classification ; yu2018efficient ; wei2020combating .
Recent work (einbinder2024label, ) investigates the noise robustness of online conformal prediction under uniform label noise, with a strong distributional assumption. Their analysis demonstrates that noisy labels will lead to a conservative long-run coverage guarantee, with the assumption that noisy score distribution stochastically dominates the clean score distribution, i.e., $\mathbb{P}\left\{\tilde{S}\leq s\right\}\leq\mathbb{P}\left\{S\leq s\right\},\forall s\in\mathbb{R}$. The distributional assumption is too strong to ensure valid coverage under general cases. Moreover, their analysis fails to quantify the specific deviation of coverage guarantees. These limitations motivate us to establish a general theoretical framework for this problem and develop a noise-robust algorithm for online conformal prediction.
3 The impact of label noise on online conformal prediction
In this section, we theoretically analyze the impacts of uniform label noise on ACI (gibbs2021adaptive, ). In this case, the threshold $\hat{\tau}_{t}$ is updated by